### v1

# PseudoNeg-MAE: Self-Supervised Point Cloud Learning using Conditional Pseudo-Negative Embeddings

Abstract
We propose PseudoNeg-MAE, a novel self-supervised learning framework that enhances global feature representation of point cloud mask autoencoder by making them both discriminative and sensitive to transformations. Traditional contrastive learning methods focus on achieving invariance, which can lead to the loss of valuable transformation-related information. In contrast, PseudoNeg-MAE explicitly models the relationship between original and transformed data points using a parametric network COPE, which learns the localized displacements caused by transformations within the latent space. However, jointly training COPE with the MAE leads to undesirable trivial solutions where COPE outputs collapse to an identity. To address this, we introduce a novel loss function incorporating pseudo-negatives, which effectively penalizes these trivial invariant solutions and promotes transformation sensitivity in the embeddings. We validate PseudoNeg-MAE on shape classification and relative pose estimation tasks, where PseudoNeg-MAE achieves state-of-the-art performance on the ModelNet40 and ScanObjectNN datasets under challenging evaluation protocols and demonstrates superior accuracy in estimating relative poses. These results show the effectiveness of PseudoNeg-MAE in learning discriminative and transformation-sensitive representations.
I Introduction
Self-Supervised Learning (SSL) has emerged as a crucial approach in point cloud analysis, allowing models to learn robust and generalizable representations from large volumes of unlabeled 3D data. With the growing accessibility of affordable 3D sensors and scanning devices, point clouds have become a widely used data format in applications such as autonomous driving, robotics, and augmented reality. SSL methods have achieved remarkable success in various downstream tasks involving point clouds, including 3D object classification [1, 2, 3, 4], 3D scene understanding [5, 6], and 3D part segmentation [1, 2, 3, 4]. Traditional SSL methods, particularly those based on contrastive learning, focus on learning invariant representations to input perturbations or transformations. However, this strict invariance to transformations can unintentionally ignore important information contained in the variations introduced by these transformations. Recent approaches [7, 8] address this limitation by introducing sensitivity to transformations, allowing models to explicitly encode the relationships between original and transformed point clouds in the latent space. Despite these advancements, the inherent design of their loss functions often leads to invariance as an inevitable outcome, making it challenging to learn representations highly sensitive to the transformations.
In this work, we propose a novel self-supervised learning framework PseudoNeg-MAE, which effectively captures transformation-sensitive features while maintaining discriminative representations. Our method introduces a new loss function that leverages the conditional generation of pseudo negatives to prevent the model from collapsing to an invariant solution. By incorporating pseudo-negative embeddings corresponding to various transformations applied to the input point cloud, we explicitly model the localized displacements in the embedding space caused by different transformations. This approach ensures that the learned features remain sensitive to transformations while avoiding trivial solutions. To achieve this, we propose COnditional Pseudo Negatives Embedding network (COPE). COPE takes the transformation parameters as inputs and outputs a weight used to linearly project the original embedding. In this way, we conditionally generate pseudo-negatives depending on the set of random transformations. With pseudo-negatives, we regularize COPE through our novel loss function, ensuring that different transformations produce distinct embeddings in the feature space.
Our contrastive framework is built on the Mask AutoEncoder (MAE) transformer framework, combining the strengths of both paradigms. We employ a shared Siamese encoder with a transformer backbone, initialized with weights from an MAE. We train this architecture using our novel contrastive method, enhancing the global representations of the MAE to be more discriminative and sensitive to transformations. The transformer is required to generate global embeddings for both original and transformed inputs that satisfy the linear projection constraint imposed by COPE. Here the use of pseudo-negatives becomes crucial to avoid the trivial solution where COPE simply produces identity weights for all transformations.
An advantage of our method is that it formulates relative pose estimation as an optimization problem in the embedding space, using transformation-sensitive features learned via the COPE. We introduce a novel inference algorithm that iteratively aligns source and target point cloud embeddings, achieving precise alignment even with large initial rotations. We validate the effectiveness of PseudoNeg-MAE through comprehensive experiments on shape classification and relative pose estimation tasks. On the ModelNet40 and ScanObjectNN datasets, our method achieves state-of-the-art performance, particularly under challenging protocols. For relative pose estimation tasks, our framework excels in accurately estimating relative rotations between point clouds, highlighting its capability to capture fine-grained transformation information.
II Related Works
Self-Supervised Learning for Point Cloud: Self-Supervised Learning (SSL) has become increasingly prevalent for leveraging unlabeled data, especially in the challenging domain of 3D point clouds. The inherent disorder and lack of predefined structure in point clouds necessitate diverse strategies for extracting meaningful representations. Among these, contrastive methods [6, 9, 1, 10] exploit correspondences across different views of point clouds to establish unsupervised pretraining frameworks using InfoNCE loss [11, 12]. In parallel, recent advancements have introduced transformer-based masked autoencoder approaches, such as PointBERT [13], which decodes discrete tokens from point patches, and PointMAE [3] along with PointM2AE [4], which directly predict the masked point patches rather than tokens. Extending this paradigm, Point2Vec [2] employs the principles of Data2vec [14] to facilitate feature space reconstruction in point clouds. Diverging from earlier SSL techniques that mainly induce biases towards invariant feature learning or local relationships for reconstruction, our approach focuses on learning both discriminative and transformation-sensitive representations, significantly enhancing the performance across multiple downstream tasks.
Sensitivity to Transformation in SSL methods: Contrastive learning  [12, 15, 16] has significantly advanced self-supervised representation learning; however, enforcing strict invariance to data augmentations can limit the expressive power of the learned features, especially in fine-grained tasks where details about specific transformations are crucial. To overcome this limitation, recent studies [17, 18, 7, 8] have explored enhancing models by incorporating sensitivity to transformations. One approach involves auxiliary tasks that require models to predict the transformations applied to the input data [19, 18, 17, 20], which helps preserve transformations-specific information within the representations and leads to less invariant feature spaces. However, these methods often lack guarantees of a consistent mapping between input transformations and changes in the latent representations [8]. Alternative approaches [21, 22], integrate loss functions that balance invariance with sensitivity to augmentations, enabling models to remain responsive to transformations while maintaining robustness. Additionally, recent methods [7, 8] have developed models that directly map transformations in the input space to transformations in the latent space, employing external predictors to modify representations accordingly. Despite these advancements, the inherent design of their loss functions often makes invariance an inevitable outcome [21, 8]. In our work, we address this issue by proposing a novel loss function that penalizes this trivial invariant feature spaces by leveraging the use of generated pseudo negatives. We demonstrate the effectiveness of incorporating an external predictor tailored to downstream tasks influenced by transformations. By introducing a specialized inference algorithm, we enhance model performance on tasks sensitive to these transformations. While previous methods largely focused on image-based classification, we extend these approaches to 3D point clouds, targeting both classification and relative pose estimation (i.e., registration).
Figure 1: Overview of the Proposed PseudoNeg-MAE Method: The method begins by processing a point cloud $\mathbf{x}_{i}$ and its transformed counterpart $\mathbf{x}_{i}^{+}$, extracting their global representations $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{+}$, respectively, using a shared transformer architecture. Simultaneously, COPE network outputs the weight $\Theta_{g}$ and $\{\Theta_{g_{r}}\}$ based on the input and random transformations $g$ and $\{g_{r}\}$. $\{\Theta_{g_{r}}\}$ are used to generate pseudo-negative embeddings ${\{\tilde{\mathbf{z}}_{i}^{g_{r}}\}}$ which help in learning features that are both discriminative and sensitive to the transformation.
III Proposed Approach - PseudoNeg-MAE
In this section, we introduce PseudoNeg-MAE, a method designed to learn feature representations that are both discriminative and sensitive to transformations. As shown in Fig. 1, we model the relationship between input transformations and displacements in the embedding space using linear projections. These projections are controlled by a network, COPE, which generates transformation-specific weights. We train the main network $f$ and COPE jointly. There is a risk that COPE could collapse into the identity matrix, producing invariant representations. To avoid this, we introduce a novel loss function that uses pseudo-negatives to regularize COPE and ensure that the projections are transformation-sensitive during training.
III-A Notations
Let $f$ be a parametric network that maps an input space $\mathcal{X}$ to the unit sphere. Here $f$ is a composition of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$, (i.e., $f=\mathcal{E}\circ\mathcal{D}$). Let $\mathcal{G}$ be a distribution of possible transformations in the input space $\mathcal{X}$ and $g$ denotes a transformation sampled from the distribution $\mathcal{G}$. For the given input point cloud $\mathbf{x}_{i}\in\mathcal{X}$, $g$ transforms it to a positive view $\mathbf{x}_{i}^{+}=g(\mathbf{x}_{i})$. The representations $\mathbf{z}$ of $\mathbf{x}_{i},\mathbf{x}_{i}^{+}$ is defined as $\mathbf{z}_{i}=f(\mathbf{x}_{i})\text{ and }\mathbf{z}_{i}^{+}=f(\mathbf{x}_{i}^{+})$. Both $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{+}$ reside within the same representation space $\mathcal{Z}$ of $f$. We introduce an operator $u_{g}$ that models how the transformation $g$ affects the feature representation $\mathbf{z}_{i}$.
III-B Problem Statement
Contrastive learning methods, such as InfoNCE [12], typically aim to learn a representation that is invariant to perturbations or transformations of the input data. This is achieved by encouraging similar instances to have similar representations in the embedding space, while dissimilar instances are pushed apart. This objective is equivalent to a combination of alignment $\mathcal{L}_{\text{align}}$ and uniformity $\mathcal{L}_{\text{unif}}$, as described in Eq. (3) below:
$\mathcal{L}_{\text{align}}=$ $\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{z}_{i}-\mathbf{z}_{i}^{+}\right\|_{2}^{2},$ (1)
$\mathcal{L}_{\text{unif}}=$ $\log\mathbb{E}_{(\mathbf{z}_{i},\mathbf{z}_{k})\sim\mathcal{Z}\times\mathcal{Z}}e^{-\left\|\mathbf{z}_{i}-\mathbf{z}_{k}\right\|_{2}^{2}/\tau},$ (2)
$\mathcal{L}_{\text{ssl}}=$ $\mathcal{L}_{\text{align}}+\mathcal{L}_{\text{unif}}.$ (3)
Specifically, $\mathcal{L}_{\text{align}}$ promotes the invariance property, while $\mathcal{L}_{\text{unif}}$ regularizes the learning process by preventing dimensional collapse, where $f(x)$ becomes constant for all $x$ [23]. Here $\tau$ is a temperature parameter.
While invariant SSL models facilitate the learning process by discarding the variations introduced by transformations, they often neglect potentially critical information these variations might carry. To address this issue, a group of self-supervised learning methods [7, 8, 21] has introduced the concept of sensitivity to transformation, enabling the model to explicitly encode the relationships between the original and augmented data in latent space.
These approaches introduce a modified alignment term, $\mathcal{L}_{\text{align}^{*}}=\frac{1}{N}\sum_{i=1}^{N}\left\|u_{g}(\mathbf{z}_{i})-\mathbf{z}_{i}^{+}\right\|_{2}^{2}$, where $u_{g}$ defines the relationship between the feature representations $\mathbf{z}_{i}$ and their augmented counterparts $\mathbf{z}i^{+}$ under the transformation $g$. When combined with the uniformity term from Eq. (2), this alignment term helps the model avoid dimensional collapse while enhancing its sensitivity to transformations.
However, a trivial solution to $\min_{f}\mathcal{L}_{\text{align}^{*}}$ is for all transformations $\forall g\in\mathcal{G}$ to collapse into the same embedding space (i.e., $u_{g}(\mathbf{z}_{i})=\mathbf{z}_{i}$), making the embeddings invariant to transformations. To avoid this, our method, PseudoNeg-MAE, introduces an additional constraint that produces localized displacements in the embedding space, ensuring that:
$u_{g_{r}}(\mathbf{z}_{i})\neq u_{g_{s}}(\mathbf{z}_{i});\quad$ $(g_{r},g_{s})\sim\mathcal{G}\times\mathcal{G}:g_{r}\neq g_{s}.$ (4)
III-C Sensitivity to Transformation through Pseudo Negatives
To effectively model the relationship $u_{g}$, we introduce COnditional Pseudo Negatives Embedding (COPE) network based on hypernetworks [24, 8]. Specifically, COPE takes transformation parameter $g$ as an input and outputs the weight $\Theta_{g}\in\mathbb{R}^{d\times d}$. This weight is used as a linear projection on $\mathbf{z}_{i}$ which learns the localized displacements due to the transformation $g$.
$\Theta_{g}=\text{COPE}{}(g);\quad\mathbf{z}_{i}^{+}=\Theta_{g}\mathbf{z}_{i}.$ (5)
In our framework, COPE is jointly optimized with the main network $f$. Here, COPE learns to approximate the effect of transformations through its output weights $\Theta_{g}$, while $f$ must learn representations $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{+}$ that satisfy this linear constraint dependent on the transformations $g$. This is a challenging task because it requires $f$ and COPE to closely cooperate.
A potential shortcut for COPE is to produce identity weights for all transformations, $\Theta_{g}=I$, effectively making the features invariant to transformations. This trivial solution undermines our goal of learning transformation-sensitive representations. To avoid this, we introduce the process of conditional generation of pseudo negatives using COPE. With pseudo negatives, we regularize COPE and prevent it from collapsing into the identity. This approach ensures that different transformations produce different projections, maintaining the sensitivity to the transformations. Specifically, we generate pseudo negatives by applying COPE to a set of randomly sampled transformations $\{g_{r}\}$ from $\mathcal{G}$, where $\{g_{r}\}$ are different from the input transformation $g$. For each $g_{r}$, we compute the corresponding weight $\Theta{g_{r}}$ with the COPE and generate the pseudo negative embedding $\mathbf{\tilde{{z}}}_{i}^{g_{r}}$:
$\mathbf{\tilde{{z}}}_{i}^{g_{r}}=\text{COPE}(g_{r})\mathbf{z}_{i}=\Theta_{g_{r}}\mathbf{z}_{i}.$ (6)
We incorporate these pseudo negatives into our novel loss function $\mathcal{L}_{\text{cope}}$ to regularize COPE.
$\mathcal{L}_{\text{cope}}=\log\left[\sum_{r=1}^{M}e^{-\left\|\Theta_{g}\mathbf{z}_{i}-\mathbf{\tilde{{z}}}_{i}^{g_{r}}\right\|_{2}^{2}/\tau}+e^{-\left\|\Theta_{g}\mathbf{z}_{i}-\mathbf{z}_{i}^{+}\right\|_{2}^{2}/\tau}\right],$ (7)
where $M$ is the number of pseudo negatives. This loss penalizes the model for mapping different transformations to similar embeddings, thus preventing the trivial solution. Consider the case of perfect alignment where $\mathbf{z}_{i}^{+}=\Theta_{g}\mathbf{z}_{i}$. Minimizing $\mathcal{L}_{\text{cope}}$ is equivalent to optimizing $\log\left[\sum_{r=1}^{M}e^{-\left\|\Theta_{g}\mathbf{z}_{i}-\mathbf{\tilde{{z}}}_{i}^{g_{r}}\right\|_{2}^{2}/\tau}+1\right]$, which effectively maximizes the pairwise distances between $\{\Theta_{g_{r}}\mathbf{z}_{i}\}$. Achieving this requires diverse $\{\Theta_{g_{r}}\}$ from the COPE, thereby penalizing trivial invariant solutions. We define the final loss $\mathcal{L}_{\text{PseudoNegMAE}}$ as follows:
$\mathcal{L}_{\text{PseudoNegMAE}}=\mathcal{L}_{\text{align}^{*}}+\beta\mathcal{L}_{\text{cope}}+(1-\beta)\mathcal{L}_{\text{unif}}.$ (8)
Specifically, $\mathcal{L}_{\text{unif}}$ encourages discriminative features and $\mathcal{L}_{\text{cope}}$ preserves the sensitivity to transformations within the feature space. This combination enables the model to learn discriminative and sensitive representations to input transformations. Fig. 2 illustrates the objective of our loss in terms of the anchor, the positive, and the pseudo-negatives. The loss $L_{\text{align}^{*}}$ encourages the positive to be pulled towards the anchor such that $\Theta_{g}$ satisfies the Eq. (5). $\mathcal{L}_{\text{cope}}$ encourages the pseudo negatives to be pushed away from the anchor. For a given $\mathbf{z}_{i}$ all the pseudo-negatives $\{\Theta_{g_{r}}\mathbf{z}_{i}\}$ only depend on the output of the COPE $\Theta_{g_{r}}$ network. This avoids $\Theta_{g}$ collapsing into identity, preventing the undesired invariant solution.
III-D COPE Network Architecture
COPE generates linear weights conditioned on the transformation set ${\{g}\}$. In our work, we consider rotations as transformations and utilize quaternions to represent them as inputs to the COPE. The network architecture is illustrated in the right section of Fig. 1.
First, we embed $g$ into a harmonic space $H(\cdot)$, using a set of high-frequency functions [25]. Subsequently, these embedded quaternions are processed through an MLP network to extract transformation-sensitive embeddings, denoted as $\mathbf{h}_{g}^{\prime}\in\mathbb{R}^{\frac{d}{k}}$. Independently estimating each element of $\Theta_{g}$ is computationally expensive. Therefore, we utilize shared learnable embedding vectors $\Psi\in\mathbb{R}^{\frac{d}{k}\times d}$, which are responsible for the entries of the rows in $\Theta_{g}$. We perform an element-wise multiplication of these vectors with our $h_{g}^{\prime}$, i.e., $\Gamma=h_{g}^{\prime}\odot\Psi$, where $\Gamma\in\mathbb{R}^{d\times\frac{d}{k}}$. A shared non-linear MLP network is then used to expand each column of $\Gamma$ to the dimension $d$. This process yields the linear weight $\Theta_{g}\in\mathbb{R}^{d\times d}$. Consequently, $\Theta_{g}$ is determined solely based on the input transformation.
We generate pseudo negatives using COPE by sampling a set of random rotations $\{{g_{r}}\}_{r=1}^{M}$. We uniformly sample quaternions representing 3D rotations to ensure coverage over the rotation space. Specifically, for each pseudo negative, we generate a random vector $\mathbf{v}_{r}=(w_{r},x_{r},y_{r},z_{r})$ by sampling from a standard normal distribution $\mathbf{v}_{r}\sim\mathcal{N}(0,I_{4})$. We then normalize $\mathbf{v}_{r}$ to obtain a unit quaternion $g_{r}$ and ensure the scalar part is nonnegative:
$g_{r}=\frac{\mathbf{v}_{r}}{|\mathbf{v}_{r}|}\cdot\text{sign}(w_{r}).$ (9)
This approach ensures that the set of quaternions ${g_{r}}$ is uniformly distributed over the space of 3D rotations. Random rotations are fed to the COPE for generating pseudo negatives as in Eq. 6.
III-E PseudoNeg-MAE Transformer Backbone
Our point cloud transformer backbone is based on a Siamese encoder-decoder network, denoted as $f$, in which there are two branches with shared weights. PseudoNeg-MAE operates on pairs of point clouds: a randomly sampled point cloud $\mathbf{x}_{i}$ and its transformed point cloud $\mathbf{x}_{i}^{+}$ with the applied transformation $g$. Following the standard transformer architecture on the point cloud, we first extract a sequence of non-overlapping point patches, which are then converted to a set of tokens. Given an input point cloud $\mathbf{x}$, we first used Furthest Point Sampling (FPS) to sample a set of $n$ center points. Then, K Nearest-Neighbour (KNN) is utilized to select $k$ neighboring points around each center point to construct $n$ point patches. Following this, patch tokens are generated using a mini-PointNet [3] and fed to the transformer encoder. However, in PseudoNeg-MAE, not all the patch tokens are processed by the encoder. A random masking strategy is applied independently to $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{+}$, masking $m$ number of the point patches. Only visible tokens are fed to an encoder ($\mathcal{E}$). This independent masking strategy encourages the learning of representations that are robust to noise and occlusions.
Then, the encoded embeddings of visible tokens from the two inputs are sent to the decoders with shared weights. An extra learnable [CLS] token is appended to each encoder output. We implement a lightweight decoder $\mathcal{D}$. The objective of the decoder is to aggregate the encoder outputs into representations $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{+}$.
Figure 2: Conceptual visualization of the loss $\mathcal{L}_{\textit{PseudoNeg-MAE}{}}$. Using COPE, we generate the anchor $\Theta_{g}\mathbf{z}_{i}$ and a set of pseudo negatives $\Theta_{g_{r}}\mathbf{z}_{i}$ for the correspondoing positive $\mathbf{z}_{i}^{+}$. As per our main goal, due to these pseudo negatives, the model remains sensitive to different transformations rather than becoming completely invariant to them.
III-F Relative Pose Estimation Using COPE
To tackle the challenge of relative pose estimation between source and target point clouds, we introduce a novel algorithm that utilizes the COPE network to optimize the transformation parameters directly. The algorithm focuses on estimating the quaternions to minimize the difference between the transformed source and target embeddings, thus achieving precise alignment.
The process, detailed in Algorithm 1, starts by exploring a set of randomly generated quaternions $\{g_{r}\}$, representing potential rotations between the source and target. Each quaternion is evaluated in a bi-directional manner—applying the transformation to both the source and the target embeddings via $g_{r}$ and $g_{r}^{-1}$. Here $g_{r}^{-1}$ is the inverse of $g_{r}$. The quaternions are refined through an iterative process, where each is adjusted according to the gradient update. After updating the quaternions, it is essential to ensure they remain valid unit quaternions, which is achieved by the normalization step at line 1. The quaternion associated with the minimum loss is designated as $g_{est}$ serving as the final solution for the alignment. This process ensures that the iterative refinement systematically approaches the global minimum of the loss landscape. In summary, this algorithm extends the capabilities of COPE to solve relative pose estimation by optimizing quaternions to align global embeddings of point clouds.
Data: Learning rate $\epsilon=0.01$
Data: Source and target embeddings $\mathbf{z}_{src},\mathbf{z}_{tgt}$
1 Initialize $\mathcal{L}_{min}=\infty$
2for j in $N_{out}$ do
3       Initialize $g_{r}=(w_{r},x_{r},y_{r},z_{r})$
4       for k in $N_{in}$ do
5             $\mathbf{z}^{g_{r}}=\mathrm{\text{COPE}{}}(g_{r})\mathbf{z}_{src}$
6             $\mathbf{z}^{g_{r}^{-1}}=\mathrm{\text{COPE}{}}(g_{r}^{-1})\mathbf{z}_{tgt}$
7             $\mathcal{L}_{g_{r}}=\left\|\mathbf{z}_{tgt}-\mathbf{z}^{g_{r}}\right\|_{2}^{2}+\left\|\mathbf{z}_{src}-\mathbf{z}^{g_{r}^{-1}}\right\|_{2}^{2}$
8             $g_{r}^{*}=g_{r}-\epsilon\nabla_{g_{r}}L_{g_{r}}$
9             $g_{r}=\frac{g_{r}}{\|g_{r}\|}\cdot\text{sign}(w_{r}^{*})$
10       end for
11
12      if $\mathcal{L}_{g_{r}}<\mathcal{L}_{min}$ then
13             Update $\mathcal{L}_{min}=\mathcal{L}_{g_{r}}$, set $g_{est}=g_{r}$
14
15       end if
16
17 end for
return $g_{est}$
Algorithm 1 Optimization of Transformation Estimation with Minimum Loss Tracking.
IV Experiments
IV-A Pre-training and implementation details
We initialize the encoder $\mathcal{E}$ of our network $f$ using a point cloud MAE [2] and pretrain with our contrastive formulation. The proposed PseudoNeg-MAE is a modular component that integrates with any point cloud MAE backbone. We follow the procedure utilized in Masked Autoencoders for images [26, 27] as an effective way of integrating the contrastive objective with mask reconstruction pre-training. We use the training set of the ShapeNet dataset [28], which contains approximately 42,000 synthetic point clouds across 55 object categories for training. We sample 1024 points from each point cloud for pretraining.
The encoder $\mathcal{E}$ comprises 12 layers, each with an internal dimension of 384. The decoder $\mathcal{D}$ is implemented as a single transformer layer with [CLS] token. Training is conducted over 1600 epochs with a batch size of 512. We utilize the AdamW optimizer, starting with a learning rate of $1\times 10^{-4}$ and applying a cosine annealing with a linear warmup scheduler. After thorough validation of the ShapeNet dataset, we set the $\beta$ parameter of the loss function to 0.3 and the number of pseudo negatives to eight.
TABLE I: Shape classification results on ModelNet40 dataset and ScanObjNN’s OBJ-BG subset under z/z and S0(3)/SO(3) evaluation scenario.
Method ModelNet40 ScanObjNN–OBJ-BG
z / z SO(3) / SO(3) SO(3) / SO(3)
Supervised
PointNet [29] 89.2 75.5 54.7
PoinNet++ [29] 89.3 85.0 47.4
PCT [30] 90.3 88.5 45.8
SFCNN [31] 91.4 90.1
RIConv [32] 86.5 86.4 78.1
RI-GCN [33] 89.5 89.5 80.6
GCAConv [34] 89.0 89.2 80.3
RI-Framework [35] 89.4 89.3 79.9
VN-DGCNN [36] 89.5 90.2 80.3
OrientedMP [37] 88.4 88.9 77.2
Yu et al. [38] 91.0 91.0 86.3
Self-Supervised Learning
PointMAE [3] 90.7 89.3 83.3
Point2Vec [2] 91.3 88.4 84.7
PseudoNeg-MAE 92.5 91.1 89.0
TABLE II: Relative pose estimation results over five categories with E2PN protocol. Here all the other compared methods are supervised. Among them, KPConv is a non-equivariant method while EPN and E2PN are equivariant networks. Our self-supervised approach performs better than the supervised equivariant methods.
Learning Type Mean / Max / Med (°) Airplane Bottle Car Chair Sofa
Supervised Learning KPConv [39] 12.0 / 70.5 / 10.0 8.0 / 104.4 / 5.0 35.8 / 175.6 / 20.9 26.0 / 168.3 / 16.3 84.2 / 177.3 / 75.2
EPN [40] 1.3 / 6.2 / 1.1 1.2 / 22.3 / 0.8 2.6 / 117.6 / 1.1 1.2 / 9.0 / 1.0 1.5 / 15.0 / 1.1
E2PN [41] 1.5 / 11.1 / 1.1 1.9 / 46.1 / 1.1 3.3 / 74.0 / 1.5 2.9 / 38.2 / 1.8 2.7 / 33.6 / 1.9
[…]
V Results on Downstream Tasks
We evaluate our model on downstream tasks, specifically shape classification and relative pose estimation for both objects and scene levels, to demonstrate the advantage of discriminative and rotation-sensitive properties of our approach.
V-A Shape Classification
We evaluate our model on two prominent 3D object classification datasets, namely, ModelNet40 and ScanObjectNN. ModelNet40 [42] consists of 12,311 clean 3D CAD models from 40 categories, split into 9,843 instances for training and 2,468 for testing. ScanObjectNN [43] is a more challenging dataset derived from real-world indoor scenes containing approximately 15,000 objects across 15 categories. It includes three variants: OBJ-ONLY with isolated segmented objects, OBJ-BG featuring objects with contextual background elements, and OBJ-T50-RS with perturbations in object extraction.
We follow the protocols of [44] for the evaluation in Tab. I: z/z – training and testing under rotations around the z-axis and SO(3)/SO(3) – training and testing the under the arbitrary 3D rotations. Tab. I provides a detailed comparison of different point cloud backbones applied to the ModelNet40 and ScanObjectNN datasets across different protocols. The table is divided into sections detailing Supervised and SSL approaches. Here supervised section consists of equivariant methods [31] - [38] which are designed to handle rotations. PseudoNeg-MAE, our SSL-based model surpasses the performance of both supervised and self-supervised methods. Specifically, in the challenging SanObjNN OBJ-BG scenario, PseudoNeg-MAE achieves 89.0%, which is a significant improvement over the best supervised method and leads other SSL methods by a considerable margin. This demonstrates that our novel pretraining approach enhances the generalization capabilities across diverse ranges.
V-B Relative Pose Estimation
We assess our model on relative pose estimation benchmark tasks, where the goal is to predict the relative rotation between pairs of point clouds. We fine-tune the network $f$ and COPE using $\mathcal{L}_{\text{align}^{*}}$ and $\mathcal{L}_{\text{cope}}$, and conduct inference following Algorithm 1. The evaluation is conducted under two scenarios: object-level and scene-level point clouds. We use mean isotropic rotation error [45] for the evaluation.
Object Level
In the object-level evaluation, we utilize five categories from the ModelNet40 dataset, following the protocol established by EPN [40]. We present both quantitative and qualitative evaluations in Tab. II and Fig. 3, respectively. Our self-supervised model, PseudoNeg-MAE, shows the best overall performance compared to supervised counterparts, including both non-equivariant (such as KPConv [39]) and equivariant (EPN [40] and E2PN  [41]) networks. By achieving lower maximum rotation errors across all categories, PseudoNeg-MAE demonstrates its superior ability to learn robust rotation-sensitive embeddings. This helps to significantly enhance the alignment accuracy of point clouds, validating the effectiveness of our proposed method in handling all ranges of rotations.
Figure 3: Illustration of relative pose estimation between source (yellow) and target (blue) point clouds with for KPConv, EPN, E2PN, and PseudoNeg-MAE. Red indicates transformed source clouds.
Scene Level
Figure 4: Relative pose estimation on 7Scenes under varying maximum rotational angles (logarithmic scale for better visualization). FMR [46], EquivReg [47], PseudoNeg-MAE.
In addition to the object-level evaluation, we extend our analysis to relative pose estimation at the scene level using the 7Scenes dataset, a real-world indoor dataset that is a subset of 3DMatch [48]. This dataset features over 100,000 points per scene. We sample 1024 points to form the source and target point clouds. We compare our method with correspondence-free pose estimation methods where FMR [46] is an iterative approach and EquivReg [47] uses an equivariant encoder. Following the protocol described EquivReg [47], we generate random rotations by uniformly sampling the angles with the specified maximum $\left(0^{\circ},\ldots,180^{\circ}\right)$, alongside randomly selected rotation axes, to compute rotation errors. The findings are summarized in Fig. 4 with a logarithmic scale for better visualization. It shows the PseudoNeg-MAE’s ability to maintain consistent lower error across the range of angles.
VI Conclusion
In this paper, we proposed PseudoNeg-MAE, a novel self-supervised learning framework aimed at preventing collapse to invariant solutions in contrastive learning for point cloud data. Our approach introduces a loss function that leverages pseudo-negatives, ensuring that the learned representations remain both discriminative and sensitive to transformations. We also presented a novel inference algorithm for relative pose estimation, which utilizes the transformation-sensitive feature space and COPE network to iteratively estimate the relative pose between point clouds. We conducted extensive experiments on shape classification and relative pose estimation tasks which demonstrate that PseudoNeg-MAE outperforms existing methods, especially in challenging rotation scenarios.